OpenStudy (buttermequeasy):

The table below shows the ticket rates for whale watching trips offered by Dolphin Tricks & Tours. Age Ticket Price Under three years free Three to twelve years $35 Over twelve years $46 On a certain day, the company took 88 people on whale watching trips. There were 44 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (44 + x)35 + (88 – (44 + x)) ∙ 46 = C (88 - x)35 + (88 – 44

OpenStudy (buttermequeasy):

Possible Answers: (44 + x)35 + (88 – (44 + x)) ∙ 46 = C (88 - x)35 + (88 – 44) ∙ 46 = C (44 – x)35 + (88 – 44) ∙ 46 = C (88 + x)35 + (88 – (44 + x)) ∙ 46 = C

OpenStudy (amistre64):

i remember that dog off of one of my kids cartoon shows :)

OpenStudy (buttermequeasy):

Aw, thanks for answering. :c

OpenStudy (amistre64):

88 people on whale watching trips. 44 children aged twelve and under, of which some children were under three years. x represents the number of children under three years which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? well, 88-44 = the number of adults at elast

OpenStudy (amistre64):

so b or c at the moment

OpenStudy (amistre64):

and c has the correct set up for 44-x(free) * price